# An approach based on metaheuristic algorithms to the timetabling problem in deregulated railway markets

### v1

@@ 3 Formulation and Simulation of the Proposal
First of all, it is necessary to detail the relationship between the IM and the different RUs. The scheduling of train timetables is a complex problem to deal with in a liberalized railway market due to the occurrence of conflicts between the different operators participating in the market. Conflicts occur due to the desire of several trains to access the track in the same time interval. These situations are handled by the IM according to the following process (Figure 1): (1) The RUs first formulate their proposed schedules to the IM, requesting the time slots and routes necessary to offer their services. (2) IM solves the timetabling problem following some criterion. In this study, the main objective is to maximize the profit of the IM. To solve the problem, this paper sets out an optimization problem which will be solved by means of metaheuristic algorithms. (3) Once the timetable has been established, it will be up to the passengers to decide which RU they want to travel with, and on which service. The process detailed here will be repeated after a certain period of time, which in the case of Spain means annually.
Given the difficulty of obtaining real data on this process, simulation is used here. Specifically, a modular system has been designed that replicates the previous process and generates the different elements necessary to carry out the tests. First, Section 3.1 presents the mathematical formulation of this optimization problem. Then, Section 3.2 shows the structure of the designed modular system, and finally, Section 3.3 describes the general outline of the optimization process.
@@ 3.1 Optimization problem
$\max\left[\sum_{i}B_{i}\cdot\left(ca_{i}-\alpha_{i}\cdot P_{{DT}_{i}}-\sum_{(j,k)\in R_{i}}\beta_{ijk}\cdot P_{{TT}_{i}}\right)\right]$ (1)
@@
  • $P_{{DT}_{{i}}}$: is the maximum penalty applied for changes to the departure time associated with service $i$.
  • $\alpha_{i}$: weights the penalty on the revenue due to deviation in the service’s departure time with respect to the proposal of the RU operating that service. This weight is obtained by the time difference between the RU’s request and the allocation made by the IM, as well as by the value of the constant $k$ for each RU to determine the RU’s sensitivity to schedule changes.
  • $R_{i}$ represents the ordered set of station pairs $(j,k)$ that are visited sequentially by service $i$. Each pair $(j,k)$ indicates a direct travel segment between stations $j$ and $k$ along the route of service $i$.
  • $P_{TT_{i}}$: is the maximum penalty applied for changes to the requested travel time between a pair of stations $(j,k)$ visited by service $i$.
  • $\beta_{ijk}$: weights the penalty on the revenue applied due to deviation in travel time between a pair of stations $(j,k)$ visited by the service $i$. Travel times shorter than the ones requested are not allowed. This weight depends on the travel time deviation between pairs of stations, as well as on the value of the constant $k$ for each RU to represent the RU’s sensitivity to schedule changes.
In summary, the objective function presented in Equation 1 reflects the IM’s goal of maximizing profit by balancing revenue from scheduled services against penalties incurred due to deviations from the requested schedules. This balance ensures that only services meeting the operational constraints are scheduled, while changes to departure times or travel times are penalized according to the sensitivity of each RU.
@@ 3.2 Modular system designed
A modular system has been developed to perform simulations and test different optimization methods, as shown in Figure 2. Each module function is now detailed.
Figure 2: Designed Modular System
@@ 3.2.1 Generation Module
This module generates a set of requests for services addressed to the IM in order to obtain the reservation of the tracks necessary to meet the timetables desired by each RU. The requests detail the timetable for the use of each service, specifically, for each request, the stations to be visited and the arrival and departure times at each of these stations are indicated. It is a module taken from the ROBIN [62] simulator which generates timetable requests in a pseudo-random way. One important aspect is the generation of the $ca_{i}$ value. As seen in Equation 1, this value relates to the willingness of the RU to pay to schedule the requested service, and the associated penalties due to any changes in the original schedule are also related to this value. In this study, $ca_{i}$ is calculated for each service based on two questions: the number of stops of the service and the RU operating it. This approach recognizes that different RUs may have varying levels of financial capacity of strategic interest, justifying the adjustment of $ca_{i}$ to reflect these variations in a reasonable manner.
@@ 3.2.2 Conflict Detection Module
Table 2: Timetables of rail services organized by station
Table 2 shows the timetables of three services on the Madrid-Barcelona line. Considering that each tuple of values represents the arrival time ($t_{a}^{STA}$) of a train at the station and the departure time ($t_{d}^{STA}$). For instance, Service three initially departs Madrid at $t_{d}^{MAD}=18:00h$, making a four-minute stop at $t_{a}^{CAL}=18:50h$ in Calatayud. Subsequently, it arrives in Lleida at $t_{a}^{LLE}=20:10h$, where it makes a new four-minute stop. Finally, it arrives in Barcelona (end of the trip) at $t_{a}^{BAR}=21:20h$.
@@
The conflict representation method presented in this paper assumes that a conflict arises when, at any time over the course of a service, there is an intersection with the safety area of another service. However, considering the motion of trains as a Uniform Rectilinear Motion (URM), it is possible to simplify the search for conflicts. Then, If the start time of service $i^{\prime}$ is later than the end time of service $i$, no conflict exists, and the check concludes quickly. In other cases, it examines all relevant stations, including shared stations and those along the route of service $i$. At each station, conflicts are evaluated based on three conditions: the gap between the departure times of the two trains at the initial station is smaller than the safety margin; the gap between their arrival times at the final station is smaller than the safety margin; and even if the time gaps are sufficient, a conflict arises if their signs differ, indicating that the paths of the two trains intersect at some point along the route. By assuming that the trains’ trajectories follow a URM, it is not necessary for service $i^{\prime}$ to visit exactly the stations of the path being checked to obtain the time value at each station; these values can be inferred from the equation of a line given two points.
@@ 3.2.3 Penalty Calculation Module
The objective of the IM, as previously mentioned, is to maximize its profit. To do so, given a series of requests that may present conflicts, the IM will generate a feasible timetable in which it will plan a series of services from all the requests received. The profit associated with the final schedule depends on the access fee that each RU pays to the IM for the use of the infrastructure. However, the schedule adjustments that the IM makes to the received requests result in penalties on this fee, reducing its profit. Therefore, the greater the deviation planned by the IM for a service’s schedule from the initially submitted proposal, the higher the penalty. This penalty on the fee affects, on the one hand, the deviation in the departure time of the service and, on the other, the sum of each deviation in departure times at all stations the service visits along its route. Suppose a train travels from Madrid to Calatayud to Zaragoza to Barcelona. The penalty associated with the fee paid by the RU for this service is related to the deviation that the IM has planned from the departure time of this train compared to the proposal received from the RU. Additionally, it also depends on the sum of the deviations in departure times at each station visited along the route, in this case, Calatayud and Zaragoza. This penalty will be different for each RU, i.e., each RU will be affected differently. It will be modeled with what has been called the “penalty sensitivity curve”.
@@
  • The parameter $P_{DT}$ is introduced as the maximum penalty for changes in the service’s departure time. In the tests reported in this study, $P_{DT}=0.35$, which means that a maximum of 35% of the penalty imposed on a service fee is due to changes in its departure time.
  • The parameter $P_{TT}$ is introduced as the penalty on the access fee associated with changes in the travel times of the service at its stations visited. This value is divided among all the origin-destination pairs of the service, so if it is a direct service (no intermediate stops), the entire penalty on the fee is due to the deviation in the train’s departure time. In this study, $P_{TT}=0.65$.
@@
To model the sensitivity curve to the penalty, Equation 3 is used.
@@
Figure 3: Penalty Sensitivity Curve
In order to calculate the revenue obtained with an IM proposal, a maximum penalty is considered for changes in the service departure times ($P_{DT}$) and changes in each one of the travel times ($P_{TT}$) for each service $i$. For example, for a maximum deviation of 10 minutes between proposals and requests, if no modification is made with respect to the requested departure time, the difference would be zero ($x=0$). In this case, the penalty associated with changes in departure time will be multiplied by this value (i.e. no penalty would be applied). On the other hand, if the proposed departure time is modified by 10 minutes with respect to that requested, the maximum departure time penalty will be applied.
@@
$x_{DT}=\frac{IM_{DT}-RU_{DT}}{interval}$ (5)
where: $IM_{DT}$ is the IM proposed departure time, $RU_{DT}$ is the RU requested departure time and $interval$: defines the maximum interval in which the requested travel times can be modified. This study considers a maximum of 10 minutes (i.e. service departure times can be moved back or forward a maximum of 10 minutes).
@@
where $-1\leq\beta_{ijk}\leq 1$.
$x_{TT}=\frac{IM_{TT}-RU_{TT}}{interval}$ (7)
where $IM_{TT}$ is the IM proposed travel time between a pair of stations, $RU_{TT}$ the RU requested travel time between a pair of stations and $interval$: defines the maximum interval in which the requested departure times can be modified (10 minutes).

### v2

@@ 3 Formulation and simulation of the proposal
The scheduling of train timetables is a complex problem to deal with in a liberalized railway market. The model of railway path request and assignment phases is part of a harmonized European approach promoted by RailNetEurope (RNE) and supported by the European Union Agency for Railways (ERA)  [^4] https://www.era.europa.eu/domains/technical-specifications-interoperability/operation-and-traffic-management-tsi_en. This model is implemented through the Path Coordination System (PCS), an international platform that enables coordination of path requests among infrastructure managers, railway undertakings, and freight corridors, ensuring consistency between requests and capacity offers. Moreover, the PCS system complies with the technical requirements established in the TAF-TSI regulation (Technical Specifications for Interoperability for Telematics Applications), which governs the interoperability of capacity and timetable management systems across the European Union. This approach is used by countries such as Spain, Switzerland, Germany, and France, among others, particularly in the context of international traffic and annual timetable planning.
Figure 1: Relationship between the RUs and IM.
Figure 1 shows how this process plays out, using a flowchart that provides a structured representation of the key phases in the railway path request and assignment process, from capacity planning to the publication of the annual timetable. This process begins approximately eleven months before the timetable is published, during the Capacity Planning phase, where early path requests (pre-request) are submitted by the RUs and tentative assignments (pre-assignment) are made by IM (Phases 1 and 2). Conflicts between requests are then evaluated, and available capacity is adjusted (Phase 3), culminating in the publication of a draft capacity plan (Phase 4). Subsequently, during the Timetable Generation phase, formal path requests (late request) are processed along with their corresponding definitive assignments (late assignment), phases 5 and 6 in the figure. Remaining conflicts are resolved (Phase 7), and ad hoc requests—those submitted outside standard deadlines—are handled (Phase 8). Finally, the definitive annual timetable is published, integrating all confirmed assignments (Phase 9). This approach enables a progressive and flexible management of railway capacity, adapting to the evolving needs of railway undertakings over time.
In this paper, the research carried out can be useful in phases 2 to 4, helping to optimize the process and resolve the conflicts generated. In this case, it focuses on the maximization of the IM benefit and the payment of penalties to the RUs affected by the changes. To solve the problem, this paper sets out an optimization problem which will be solved by means of metaheuristic algorithms. Given the difficulty of obtaining real data on this process, simulation is used here. Specifically, a modular system has been designed that replicates the previous process and generates the different elements necessary to carry out the tests. First, Section 3.1 sets out the mathematical formulation of this optimization problem. Then, Section 3.2 shows the structure of the modular system designed, and finally, Section 3.3 describes the general outline of the optimization process.
@@ 3.1 Optimization problem
${\scriptstyle\max\left[\sum_{i=1}^{|SS|}B_{i}\cdot\left(ca_{i}\cdot\left(1-\alpha_{i}\cdot P_{DT_{i}}-\sum_{(j,k)\in R_{ijk}}\beta_{ijk}\cdot\frac{P_{TT_{i}}}{|R_{ijk}|}\right)\right)\right]}$ (1)
@@
  • $P_{DT_{i}}$: is the maximum penalty applied for changes to the departure time for service $i$.
  • $\alpha_{i}$: weights the penalty on the revenue due to deviation in the service’s departure time with respect to the proposal of the RU operating that service. This weight is obtained by the time difference between the RU’s request and the allocation made by the IM, as well as by the value of a constant for each RU to determine the RU’s sensitivity to schedule changes.
  • $R_{ijk}$ represents the ordered set of station pairs $(j,k)$ that are visited sequentially by service $i$. Each pair $(j,k)$ indicates a direct travel segment between stations $j$ and $k$ along the route of service $i$. $|R_{ijk}|$ is the size of the set (pairs of stations of the service $i$).
  • $P_{TT_{i}}$: is the maximum penalty applied for changes to the requested travel time between each pair of stations of the service $i$. The penalty percentage associated to each pair of stations of a service is obtained by dividing by the number of pairs $|R_{ijk}|$. Like $P_{DT}$, this value is also shared between all services.
  • $\beta_{ijk}$: weights the penalty on the revenue applied due to deviation in travel time between a pair of stations $(j,k)$ visited by the service $i$. Travel times shorter than those requested are not allowed. This weight depends on the travel time deviation between pairs of stations, as well as on the value of the constant $k$ for each RU to represent the RU’s sensitivity to schedule changes.
  • $SS$: set of services, where $|SS|$ is the number of services.
In summary, the objective function presented in Equation 1 reflects the IM’s goal of maximizing profit by balancing revenue from scheduled services against penalties incurred due to deviations from the requested schedules. When the RUs request their services from the IM, these services have an associated amount that the RUs are willing to pay for each service, provided that the service is planned with the same conditions as requested. In the modeling of the problem, it has been considered that any deviation from the original request will incur a penalty in the cost that the RU is willing to pay for the service. Therefore, this is a decrease in the benefit to the IM when it makes changes to the requests. The more variation, the greater the penalty. This model allows incorporating the case without penalties by setting these to zero by means of the variables $P_{{DT}_{{i}}}$ and $P_{TT_{ijk}}$. This balance ensures that only services meeting the operational constraints are scheduled, while changes to departure times or travel times are penalized according to the sensitivity of each RU.
@@
3.1.1 Problem constraints
This section provides the formal definition of the problem constraints: departure time interval, operational times, dwell times and conflicts between services.
Departure time interval: The modified departure times ($\tilde{d}t_{i}^{1}$) by the IM must not exceed a specified margin $\delta$, either behind or in advance of the requested ones.
$dt_{i}^{1}-\delta\leq\tilde{d}t_{i}^{1}\leq dt_{i}^{1}+\delta$ (2)
where:
  • $dt_{i}^{j}$ is the original departure time at station $j$ of service $i$. In this first equation, only the departure time at the first station is considered.
  • $\tilde{d}t_{i}^{j}$ is the updated departure time at the station $j$.
  • $\delta$ is the IM modification margin.
Operating time constraint: the departure time proposed by the IM for each station ($\tilde{d}t_{i}^{j}$) must be greater than or equal to the departure time of the previous station ($\tilde{d}t_{i}^{j-1}$) plus the original operating times (travel and dwell times). This constraint applies from the second until the penultimate stations. The first and last are excluded since the first lacks a previous station, and the last would not a have a departure time.
$\tilde{d}t_{i}^{j}\geq\tilde{d}t_{i}^{j-1}+\tau_{i}^{j-1}+\sigma_{i}^{j}\quad(j=2,\dots,|R_{ijk}|-1)$ (3)
where:
  • $\tau_{i}^{j-1}$ is the travel time from stop $j-1$ to stop $j$ for service $i$.
  • $\sigma_{i}^{j}$ is the minimal dwell time at stop $j-1$ for service $i$.
Dwell time constraints: the departure time from station $j$ must satisfy the following two restrictions:
  1. 1.
    The departure time of station $j$ ($\tilde{d}t_{i}^{j}$) must be less than or equal to the sum of the original departure time ($dt_{i}^{j}$) and maximum dwell time extension ($\Delta_{\max}$).
    $\tilde{d}t_{i}^{j}\leq\,dt_{i}^{j}+\Delta_{\max}\quad(i=2,\dots,|R_{ijk}|-1)$ (4)
  2. 2.
    The departure time from station $j$ ($\tilde{d}t_{i}^{j}$) must be less than or equal to the sum of the departure time from the previous station ($\tilde{d}t_{i}^{j-1}$) plus operational travel time ($\tau_{i}^{j-1}$) plus maximum dwell time extension ($\Delta_{\max}$).
    $\tilde{d}t_{i}^{j}\leq\tilde{d}t_{i}^{j-1}+\tau_{i}^{j-1}+\Delta_{\max}\quad(i=2,\dots,|R_{ijk}|-1)$ (5)
where:
  • $at_{i}^{j}$ is the original arrival time at station $j$ of service $i$.
  • $\Delta_{\max}$ is the maximum allowable dwell-time.
Conflicts between services: In order to identify any conflict between service $i$ and service $i^{\prime}$, the values $\Delta dt_{i,i^{\prime}}^{j}$ and $\Delta at_{i,i^{\prime}}^{j+1}$ are calculated in the first instance (Equation 6 and 7). This values represent the time gap of the departure and arrival times between a common pair of stations for two services:
$\Delta dt_{i,i^{\prime}}^{j}\;=\;\tilde{d}t_{i}^{j}-\tilde{d}t_{i^{\prime}}^{j}$ (6)
where $\Delta dt_{i,i^{\prime}}^{j}$ is the departure‐time gap between service $i$ and $i^{\prime}$ at station $j$.
$\Delta at_{i,i^{\prime}}^{j+1}\;=\;\tilde{a}t_{i}^{j+1}-\tilde{a}t_{i^{\prime}}^{j+1}$ (7)
where $\Delta at_{i,i^{\prime}}^{j+1}$ is the arrival‐time gap between service $i$ and $i^{\prime}$ at station $j+1$.
A conflict between two services arises when: the gaps have opposite signs (meaning that the two services intersect), the departure or arrival gap is less than twice the safety margin (because it is considered that each train carries its own safety margin). For example, considering two services $i$ and $i^{\prime}$ with the following schedules between two shared stations:
  • Service $i$: $\tilde{d}t_{i}^{j}=0$ and $\tilde{a}t_{i}^{j+1}=40$.
  • Service $i^{\prime}$: $\tilde{d}t_{i^{\prime}}^{j}=5$ and $\tilde{a}t_{i^{\prime}}^{j+1}=35$.
Applying Equations 6 and 7 with the previous schedule times, the values $\Delta dt_{i,i^{\prime}}^{j}\;=\;\tilde{d}t_{i}^{j}-\tilde{d}t_{i^{\prime}}^{j}=0-5=-5$ and $\Delta at_{i,i^{\prime}}^{j+1}\;=\;\tilde{a}t_{i}^{j+1}-\tilde{a}t_{i^{\prime}}^{j+1}=40-35=5$ are obtained. This implies an overtaking of service $i^{\prime}$ to $i$ (service $i^{\prime}$ departed after service $i$ but arrived earlier), which is not possible when considering a single track.
Formally, this constraint is represented as in Equation 8.
$C_{i,i^{\prime}}\;=\;\begin{cases}1,&\begin{aligned} &\exists\,(j,k)\in R_{ijk}\,\cap\,R_{i^{\prime}jk},\;B_{i}=B_{i^{\prime}}=1,\;\\ &\quad\Delta dt_{i,i^{\prime}}^{\,j}\ \cdot\Delta at_{i,i^{\prime}}^{\,j+1}\,\leq\,0,\\ &\quad\bigl{|}\Delta dt_{i,i^{\prime}}^{\,j}\bigr{|}<2\ \cdot\omega,\\ &\quad\bigl{|}\Delta at_{i,i^{\prime}}^{\,j+1}\bigr{|}<2\ \cdot\omega;\end{aligned}\\[8.61108pt] 0,&\text{otherwise.}\end{cases}$ (8)
where:
  • $C_{i,i^{\prime}}$ is the conflict indicator between services $i$ and $i^{\prime}$ ($1=$ conflict, $0=$ no conflict).
@@ 3.2 Modular system designed
A modular system has been developed to perform simulations and test different optimization methods, as shown in the flowchart in Figure 2. It is important to highlight that the diagram names the structures, although they are formally defined in Section 3.3. The first step consists of obtaining the schedule proposals from the RUs. This is done using a module of the ROBIN simulator that generates the structure $SS$ (Set of Schedules), which contains the proposals from each RU (Section 3.2.1). Next, the iterations of the metaheuristic algorithm begin, where Steps 2 to 5 occur in each iteration. $SS$ is transformed into the departure time vector $DT$, suitable as input for the metaheuristic algorithm, which produces the optimized departure times $ODT$ as output. $ODT$ may lead to conflicts between services, which are resolved in Step 4, resulting in a conflict-free vector $ODT$ and the vector $B$ from Equation 1 — the general equation of the formulation — which indicates the services that would be scheduled. With these values, the revenue is calculated by applying the penalties based on the changes made in Step 4. Then, it is checked whether a new iteration is necessary. Once all iterations have been completed, the final service assignment is obtained in $FSS$.
What follows details how the generation of schedule requests (Step 1, Section 3.2.1), conflict detection (Step 4, Section 3.2.2), and penalty calculation (Step 5, Section 3.2.3) are carried out.
Figure 2: Flowchart of the designed modular system.
@@ 3.2.1 Generation module
This module generates a set of requests for services addressed to the IM in order to obtain the reservation of the tracks necessary to meet the timetables desired by each RU. The requests detail the timetable for the use of each service, specifically, for each request, the stations to be visited and the arrival and departure times at each of these stations are indicated. It is a module taken from the ROBIN delCastillo24 simulator which generates timetable requests in a pseudo-random way. It’s considered that each RU submits only as many service requests as its allocated infrastructure can support—namely, the total number of trains that its assigned storage and maintenance yards can accommodate. In other words, the model does not permit an unbounded number of services: the input request set for each RU has been pre‐filtered so as not to exceed its yard capacity. One important aspect is the generation of the $ca_{i}$ value. As seen in Equation 1, this value relates to the willingness of the RU to pay to schedule the requested service, and the associated penalties due to any changes in the original schedule are also related to this value. In this study, $ca_{i}$ is calculated for each service based on two questions: the number of stopsthe service makes and the RU operating it. This approach recognizes that different RUs may have varying levels of financial capacity or strategic interest, justifying the adjustment of $ca_{i}$ to reflect these variations in a reasonable manner.
@@ 3.2.2 Conflict detection module
Table 2: Timetables of rail services organized by station.
@@
Table 2 shows the timetables of three services on the Madrid-Barcelona line. Considering that each tuple of values represents the arrival time ($t_{a}^{STA}$) of a train at the station and the departure time ($t_{d}^{STA}$). For instance, Service three initially departs Madrid at $t_{d}^{MAD}=18:00h$, making a four-minute stop at $t_{a}^{CAL}=18:50h$ in Calatayud. Subsequently, it arrives in Lleida at $t_{a}^{LLE}=20:10h$, where it makes a new four-minute stop. Finally, it arrives in Barcelona (end of the trip) at $t_{a}^{BAR}=21:20h$. Figure 3 shows the corresponding Marey chart for the three services presented in Table 2. Conflicts between services are highlighted in red, considering a 10 minute safety headway for each service.
Figure 3: Marey chart representation of the timetables.
@@
The conflict representation method presented in this paper assumes that a conflict arises when, at any time over the course of a service, there is an intersection with the safety area of another service. However, considering the motion of trains as a Uniform Rectilinear Motion (URM), it is possible to simplify the search for conflicts. At the strategic planning stage—i.e., the process of developing long‐range, year-ahead timetables and resource allocations several months before service deployment—a URM approximation in which each train’s between-stop trajectory is modeled as a constant-speed segment has been considered. This simplification reduces model complexity and enables efficient, network-wide conflict detection. To absorb unmodeled variability in running speeds, acceleration/deceleration profiles and station dwell times, a fixed safety headway $\omega$ has been embedded into the conflict-gap constraints—thereby reserving extra time-slots around each passage and preventing spurious conflicts even when actual operations deviate from the nominal URM profile. Then, if the start time of service $i^{\prime}$ is later than the end time of service $i$, no conflict exists, and the check concludes quickly. In other cases, it examines all relevant stations, including shared stations and those along the route of service $i$. At each station, conflicts are evaluated based on three conditions: the gap between the departure times of the two trains at the initial station is smaller than the safety margin; the gap between their arrival times at the final station is smaller than the safety margin; and even whether the time gaps are sufficient, a conflict arises if their signs differ, indicating that the paths of the two trains intersect at some point along the route. By assuming that the trains’ trajectories follow a URM, it is not necessary for service $i^{\prime}$ to visit exactly the stations of the path being checked to obtain the time value at each station; these values can be inferred from the equation of a line given two points.
@@ 3.2.3 Penalty calculation module
The objective of the IM, as previously mentioned, is to maximize its profit. To do so, given a series of requests that may present conflicts, the IM will generate a feasible timetable in which it will plan a series of services from all the requests received. The profit associated with the final schedule depends on the access fee that each RU pays to the IM for the use of the infrastructure. However, the schedule adjustments that the IM makes to the received requests result in penalties on this fee, reducing its profit. Therefore, the greater the deviation planned by the IM for a service’s schedule from the initially submitted proposal, the higher the penalty. This penalty on the fee affects, on the one hand, the deviation in the departure time of the service and, on the other, the sum of each deviation in departure times at all stations the service visits along its route. The timetabling problem addressed in this study is an annual, in-advance planning exercise, carried out several months before actual service deployment. Consequently, all market conditions used to calibrate penalty coefficients are those forecast by the respective RUs at the start of the planning cycle, and no subsequent updates to these estimates occur during the optimization. In other words, slot allocations—and hence penalty values—are based on the static, RU-provided forecasts, so time-dependent variation in penalty parameters has not been considered. Suppose a train travels from Madrid to Calatayud to Zaragoza to Barcelona. The penalty associated with the fee paid by the RU for this service is related to the deviation that the IM has planned from the departure time of this train compared to the proposal received from the RU. Additionally, it also depends on the sum of the deviations in departure times at each station visited along the route, in this case, Calatayud and Zaragoza. This penalty will be different for each RU, i.e., each RU will be affected differently. It will be modeled with what has been called the “penalty sensitivity curve”.
@@
  • The parameter $P_{DT_{i}}$ is introduced as the maximum penalty for changes in the service’s departure time. In the tests reported in this study, $P_{DT_{i}}=0.35$, which means that a maximum of 35% of the penalty imposed on a service fee is due to changes in its departure time.
  • The parameter $P_{TT_{i}}$ is introduced as the penalty on the access fee associated with changes in the travel times of the service at its stations visited. This value is divided among all the origin-destination pairs of the service, so if it is a direct service (no intermediate stops), the entire penalty on the fee is due to the deviation in the train’s departure time. In this study, $P_{TT_{i}}=0.65$.
@@
The values of $P_{max}$, $P_{DT_{i}}$, $P_{TT_{i}}$ are configurable in the model and have been adjusted empirically.
To model the sensitivity to the penalty, Equation 10 is used. The selected penalty curve tries to model how modifications on the requested times will affect the willingness to pay, considering that the modifications proposed by the IM are always within a margin (e.g. 60 minutes). If there are no changes to the request, no penalties are incurred. On the other hand, as the modification proposed by the IM approaches the maximum margin, the corresponding maximum penalty is incurred. In addition, the curve tries to model different types of sensitivity based on the value of parameter k, so that the willingness to pay of the RUs is not always affected in the same way.
@@
Figure 4: Penalty sensitivity curve.
In order to calculate the revenue obtained with an IM proposal, a maximum penalty is considered for changes in the service departure times ($P_{DT_{i}}$) and changes in each one of the travel times ($P_{TT_{ijk}}$) for each service $i$. For example, for a maximum deviation of 10 minutes between proposals and requests, if no modification is made with respect to the requested departure time, the difference would be zero ($x=0$). In this case, the penalty associated with changes in departure time will be multiplied by this value (i.e. no penalty would be applied). On the other hand, if the proposed departure time is modified by 10 minutes with respect to that requested, the maximum departure time penalty will be applied.
@@
$x_{DT}=\frac{IM_{DT}-RU_{DT}}{\delta}$ (12)
where $IM_{DT}$ is the IM proposed departure time, $RU_{DT}$ is the RU requested departure time and $\delta$ defines the maximum interval in which the requested travel times can be modified. This study considers a maximum of 10 minutes (i.e. service departure times can be moved back or forward a maximum of 10 minutes).
@@
where $0\leq\beta_{ijk}\leq 1$.
$x_{TT}=\frac{IM_{TT}-RU_{TT}}{\delta}$ (14)
where $IM_{TT}$ is the IM proposed travel time between a pair of stations, $RU_{TT}$ the RU requested travel time between a pair of stations and $\delta$ defines the maximum interval in which the requested departure times can be modified (10 minutes).